# Superposition Yields Robust Neural Scaling

### v2

@@ Appendix B Toy model training
In this Appendix, we explain how we trained the toy models and obtained raw data. There are two classes of toy models trained. The first one is a large toy model with data dimension $n=10240$, which is reported in Figure 1 and Figure 4 to show scaling behavior across around two orders of magnitude. The other toy model class is small toy models fixing $n=1000$, such that we can scan more hyperparameters. Figures 3, 4, 5, 6, and 7 use small toy models.
@@ B.1 Large toy models
  • Weight decay: $-1.0$ for strong superposition, and $0.1$ for weak superposition
@@ B.2 Small toy models
In the small toy models reported in Figures 3, 4, 5, and 6, we set the hyperparameters as
@@
  • Hidden dimension $m$: Varied logarithmically between 10 and 100, across 6 distinct sizes.
@@ C.1 Overlap analysis
[…]
@@
[…]
@@ D.1 Figure 1
The LLM data are copied from Figure 8b, with slope $-0.91\pm 0.04$ being close to $1$ as well. We will explain details about Figure 8 later.
@@
We also output the weight matrix $W$ for these large toy models (Figure 9). They follow the same pattern that in the weak superposition regime, row norms are bimodal and are either close to 00 or $1$, making $0.5$ a good separation point for measuring how many features are represented. And in the strong superposition regime, the row norms are distributed near $1$, and $1$ is a good separation point for the two peaks, i.e., the peak greater than 1 refers to strongly represented features which are more important, and the peak smaller than 1 corresponds to the weakly represented.
@@
We can analyze the large toy model in the same way as what has been done in Figure 5 and 6. The fraction of represented features $\phi_{1/2}$ is calculated, which is 1 in the strong superposition regime, while it is close to $m/n$ in the weak superposition regime (Figure 10a).
With the measured $\phi_{1/2}$, we can estimate the loss due to unlearned features, $\langle v^{2}\rangle\sum_{i=\phi_{1/2}n}^{n}p_{i}$. This loss by unlearned features agrees well with the actual loss in the weak superposition regime (Figure 10b).
@@ D.3 Figure 3
In Figure 3, we reported results from the trained small toy models with data dimension $n=1000$, whose details are in Appendix B.2.
We showed results at data exponent $\alpha=1$ in Figure 3a. The inner panels are obtained at $m=100$. We showed that the more frequent features tend to have larger norms or to be better represented. And the norm distribution is very bimodal. We here show that it is true that the norm is around $1$ or 00 for various $\alpha$ and model sizes $m$ and degrees of superposition (Figures 11 and 12). The fraction of represented, $\phi_{1/2}$, can be calculated directly.
Figure 11: Row norm distribution at weak superposition ($\gamma=0.55$) shows that the rows either are close to zero or have norm close to 1, making $0.5$ a good separation. The 17 rows of panels from top to down correspond to 17 $\alpha$ from 0 to 2. And the 6 columns from left to right correspond to 6 $m$ from $10$ to $100$.
Figure 12: Row norm distribution at strong superposition ($\gamma=-0.55$) shows that the rows have norm close to 1. And density at $1$ is very low, making $1$ a good separation point for two groups of row norms. The 17 rows of panels from top to down correspond to 17 $\alpha$ from 0 to 2. And the 6 columns from left to right correspond to 6 $m$ from $10$ to $100$.
We next showed the $\phi_{1/2}$ as a function of $\alpha$ and $\gamma$ at $m=39$. Here, we provide the same plot but at different $m$ (Figure 13). The pattern is robust, suggesting weight decay is a good tool to change the degree of superposition.
Figure 13: Fraction of represented features as a function of $\gamma$ (x-axis) and $\alpha$ (y-axis). The 6 columns from left to right correspond to 6 $m$ from $10$ to $100$. The colorbar is the same as the main text (Figure 3b) where purple means $1$.
D.4 Figure 4
Figure 4a reports the $R^{2}$ values from the fitting introduced in Appendix D.1, where the raw data comes from training large toy models (Appendix B.1). When we are fitting the loss in log-log as a line, we choose the linear part to fit. If the loss versus model dimension curve is obviously not a line, we fit the whole curve as a line and output the $R^{2}$ value as a measure of how non-linear it is. Specifically, we fit the last five points for the power-law decay feature case in the weak superposition regime (yellow data in Figure 1b). Other cases in the weak superposition regime are fitted to a line with all data. In the strong superposition regime, when feature frequency decreases as a power law or as a linear function, we fit the data as a line starting from the third point (yellow and green in Figure 1d).
@@ D.5 Figure 5
In Figure 5a, we plot the raw losses from small toy model experiments (Appendix B.2). The “loss by unlearned" is approximated by the integral, which is
$L=\left\{\begin{aligned} &\frac{\phi_{1/2}^{1-\alpha}-1}{1-n^{1-\alpha}}n^{1-\alpha},~{}\alpha\neq 1,\\ &-\frac{\ln\phi_{1/2}}{\ln n},~{}\alpha=1.\end{aligned}\right.$
@@ D.6 Figure 6
Figure 6 studies the results from small toy models (Appendix B.2) focusing on the strong superposition regime. For the strongly represented fraction, $\phi_{1}$, we can directly compute based on the definition and the obtained weight matrices (Figure 6, a, b, and c). Here, we provide more data to show that $1$ is a natural separation point in norm to determine which are strongly represented and which are weakly represented (Figure 12).
Figure 15: The overlap variance of the strongly represented is smaller than that of random vectors. We calculate the absolute overlaps of the strongly represented and can calculate the variance. We plot variance versus model dimension for different $\alpha$ (denoted by color) and weight decays (all the negative weight decays). The dashed line is obtained by sampling $m^{2}/2$ Gaussian random vectors, normalizing them to have norm 1, calculating the absolute overlaps, and calculating the variance of the overlaps.
@@
Once select the rows with norm greater than 1, we can calculate their mean square overlap based on normalized rows $W_{i}/\|W_{i}\|_{2}$ (Figure 6d). We argue that after training, the vectors will be more similar to ETFs than to random initialization. This is studied via the variance of overlaps. ETFs, in theory, have zero variance. We find that the majority of the overlap variances are much smaller than the random initialization, especially when features have similar frequencies, which agrees with the expectation (Figure 15). The cases where the actual variance is greater than that of the random vectors have large $\alpha$, roughly correspond to the cases where $\alpha_{m}$ deviates from $1$ — ETF-like configuration no longer dominates. This is intuitive that when $\alpha$ is too large, the heterogeneity of overlaps will become large — it is better to let more frequent features occupy larger angle space. We argue that the large variance at large $\alpha$ does not mean the configuration tends to be random, but tends to be something more closely related to the frequency distribution of the features.
@@
As a related side note, our explanations based on the strongly and weakly represented features capture the basic trend that when $\alpha$ is getting large, the more important features will have larger angle space and the loss decay will be more related to the data exponent. However, this theory is oversimplified, where the strongly represented all have small overlaps and the weakly represented all have large overlaps. The real situation may be more like the angle occupied by one feature decreases continuously as the frequency decreases. As suggested by Figure 15, overlap variance within the strongly represented is greater when $\alpha$ is larger. To be more precise about the overlap distribution as well as the exponent $\alpha_{m}$ when $\alpha$ is large, we cannot use simple theoretical expectations like ETFs but have to solve the toy model.
Figure 16: The weakly represented features can have large overlaps as expected, where we plotted all the data when $\gamma<0$, and the dashed line is $1/m$. The colorbar corresponds to $\alpha$ in the same way as the main text.
We also provide evidence that overlaps of the weakly represented (Figure 16). We see that all the mean square overlaps are larger than $1/m$ instead of being on the line $1/m$. And some values are around $1$ and not decreasing much with increasing model size $m$, supporting our picture that the less frequent features are squeezed into a region and have large overlaps.
@@
Figure 17: Small toy models with $m$ from $50$ to $150$ (such that $m^{2}/2>n$) in the strong superposition regime yield similar $\alpha_{m}$ around 1 when $\alpha$ is small and a slightly smaller $\alpha_{m}$ (smaller than $2(\alpha-1)$) when $\alpha$ is large. We copied Figure 6e and made the points transparent for comparison. The non-transparent points are from small toy models with $m$ from $50$ to $150$. $\alpha_{m}=1$ is the horizontal line, $\alpha_{m}=2(\alpha-1)$ is the fast increasing dashed line, and $\alpha_{m}=\alpha-1$ is the slowly increasing dashed line. Error bars are standard errors.
@@ D.8 Figure 8
Figure 20: LLMs are in strong superposition based on the non-zero norms of the representation vectors. (a) OPT models. (b) Qwen2.5 models. (c) GPT-2 models. (d) Pythia models. The dots with error bars are mean values, and the error bar is the standard deviation. Red small points are the maximum values, and dark small points are the minimum values. The mean norm value as a characteristic row norm does not depend on model size much as expected (Appendix A.2).
@@
$L=C_{m}/m^{\alpha_{m}}+L_{\backslash m},$
where $C_{m}/m^{\alpha_{m}}$ is universal and $L_{\backslash m}$ is a constant depending on the dataset and model class. There are in total $16$ different $L_{\backslash m}$ since we have $4$ different model classes and 4 datasets. In our fitting model, there are in total $18$ parameters. We use Adam to minimize the mean square error between the predicted loss by the above function and the real loss. All losses obtained are used in optimization. The code is in nonlinearfit-3.ipynb.
We provide the raw data, losses, as a function of $1/m$ (Figure 21). The losses looks like a line with $1/m$ in one model class and with the same dataset. And the slope of the line seems to be universal. These two points support us in proposing the formula above, where $C_{m}/m^{\alpha_{m}}$ is universal. The intersections are different depending on the dataset and model class, corresponding to different $L_{\backslash m}$.
@@
We study the relationship between model dimension $m$ and model size $N$ (number of parameters). For the four open-sourced models we analyzed [38, 39, 40, 41], we can see that $N\sim m^{3}$, especially when $m$ is large. If we fit the $N$-$m$ relation by a power law while assuming a universal exponent but different coefficients depending on the model class, we obtain an exponent of $2.51$ (Figure 23a). For the Chinchilla model reported in [3], we find $N$ is also close to a power law with the model dimension, and the fitted exponent is $2.52\pm 0.03$ (Figure 23b).

### v3

@@ Appendix B Toy model training
In this Appendix, we explain how we trained the toy models and obtained raw data. There are two classes of toy models trained. The first one is a large toy model with data dimension $n=10240$, which is reported in Figure 1 and Figure 9 to show scaling behavior across around two orders of magnitude. The other toy model class is small toy models fixing $n=1000$, such that we can scan more hyperparameters. Figures 3, 4, 5, 14, and 9 use small toy models.
@@ B.1 Large toy models
  • Weight decay: 1.0-1.0 for strong superposition, and $0.1$ for weak superposition
@@ B.2 Small toy models
In the small toy models reported in Figures 3, 4, 5, 14, and 9, we set the hyperparameters as
@@
  • Hidden dimension $m$: Varied logarithmically between 10 and 100 , across 6 distinct sizes, i.e., $m=10,15,25,39,63,100$.
@@ C.1 Overlap analysis
    $\text{mean\_overlap}=\frac{\sum_{i<j}\text{overlap}(W_{i},W_{j})}{n(n-1)/2}$
@@
    $\text{variance\_overlap}=\frac{\sum_{i<j}(\text{overlap}(W_{i},W_{j})-\text{mean\_overlap})^{2}}{n(n-1)/2}$
From these values, we can calculate mean square overlaps as $\text{mean\_overlap}^{2}+\text{variance\_overlap}$.
@@ D.1 Figure 1
The LLM data are copied from Figure 6b, with slope $0.91\pm 0.04$ being close to $1$ as well. We will explain details about Figure 6 later.
@@
We also output the weight matrix $W$ for these large toy models (Figure 7). They follow the same pattern that in the weak superposition regime, row norms are bimodal and are either close to 0 or $1$, making $0.5$ a good separation point for measuring how many features are represented. And in the strong superposition regime, the row norms are distributed near $1$, and $1$ is a good separation point for the two peaks, i.e., the peak greater than 1 refers to strongly represented features which are more important, and the peak smaller than 1 corresponds to the weakly represented.
@@
Figure 9: Rich scaling phenomena arise when we change the degree of superposition and data structures. (a) In the same experiments ($n=10240$) in Figure 1, where $p_{i}\propto\exp(-i/400)$ (blue), $p_{i}\propto 1/i^{1.2}$ (yellow), and $p_{i}\propto n-i$ (green), we use $\gamma=-1$ to reach strong superposition and $\gamma=0.1$ for weak superposition. R-squared value of the fitting is used to measure how likely the fitted part is a power law (Figure 10). R-squared values closer to 1 mean that the data are more similar to a power law. We found that at strong superposition, power laws are robust across different underlying feature distributions. Yet, at weak superposition, only power-law feature frequencies can lead to power-law losses. (b) When changing superposition by weight decay and varying feature frequency decay by $\alpha$ given $p_{i}\propto 1/i^{\alpha}$, we found roughly three distinct behaviors. For (b), $n=1000$ and $m=10,15,25,39,63,100$.
We can analyze the large toy model in the same way as what has been done in Figure 4 and 5. The fraction of represented features $\phi_{1/2}$ is calculated, which is 1 in the strong superposition regime, while it is close to $m/n$ in the weak superposition regime (Figure 8a).
With the measured $\phi_{1/2}$, we can estimate the loss due to unlearned features, $\langle v^{2}\rangle\sum_{i=\phi_{1/2}n}^{n}p_{i}$. This theoretical value agrees well with the actual loss in the weak superposition regime (Figure 8b).
@@
In Figure 1, we set weight decay $\gamma=-1$ to have strong superposition and $\gamma=0.1$ to have weak superposition. We compute $R$ squared values from linear fits in log-log plots to quantify scaling behavior, assessing how closely the loss follows a power law to model dimension. We can see that at strong superposition, the losses are close to power laws, regardless of the underlying feature frequencies, yet the loss is a power law at weak superposition if the feature frequency $p_{i}$ is a power law with rank $i$ (Figure 9a). For a systematic scan, we next set $p_{i}\propto 1/i^{\alpha}$ and can vary the data exponent $\alpha$ to change how fast $p_{i}$ decays consistently. Assuming a power-law form for the final test loss, $L\propto 1/m^{\alpha_{m}}$, we extract the model exponent $\alpha_{m}$ from the empirical fit. We fit the loss with a power law in all cases. The fitted $\alpha_{m}$ reveals how fast losses decay, even in the regime where the loss should not be a power law. Roughly, three distinct patterns emerge: (1) under weak superposition (positive $\gamma$, yellow box in Figure 9b), $\alpha_{m}$ is small, indicating slow loss decay; (2) under strong superposition and a wide range of small data exponents (red box), $\alpha_{m}$ remains robustly near 1; (3) for strong superposition with large data exponents (blue box), $\alpha_{m}$ increases with $\alpha$. By interpreting these three patterns, we aim to understand when loss follows a power law with model dimension, and what determines the exponent when it does.
Figure 9a reports the $R^{2}$ values from the fitting, where the raw data comes from training large toy models (Appendix B.1). When we are fitting the loss in log-log as a line, we choose the linear part to fit. If the loss versus model dimension curve is obviously not a line, we fit the whole curve as a line and output the $R^{2}$ value as a measure of how non-linear it is. Specifically, we fit the last five points for the power-law decay feature case in the weak superposition regime (yellow data in Figure 1b). Other cases in the weak superposition regime are fitted to a line with all data. In the strong superposition regime, when feature frequency decreases as a power law or as a linear function, we fit the data as a line starting from the third point (yellow and green in Figure 1d).
@@ D.3 Figure 3
In Figure 3, we reported results from the trained small toy models with data dimension $n=1000$, whose detailed hyperparameters are in Appendix B.2.
We showed results at data exponent $\alpha=1$ in Figure 3. The results are obtained at $m=100$, $\gamma=-1$ for panel a, and at different $m$ and $\gamma$ for panel b. We showed that the more frequent features tend to have larger norms or to be better represented. And the norm distribution is very bimodal. We here show that it is true that the norm is around $1$ or 0 for various $\alpha$ and model sizes $m$ and degrees of superposition (Figures 11 and 12). The fraction of represented, $\phi_{1/2}$, can be calculated directly.
Figure 11: Row norm distribution at weak superposition ($\gamma=0.55$) shows that the rows either are close to zero or have norm close to 1, making $0.5$ a good separation. The 17 rows of panels from top to down correspond to 17 $\alpha$ from 0 to 2. And the 6 columns from left to right correspond to $m=10,15,25,39,63,100$.
Figure 12: Row norm distribution at strong superposition ($\gamma=-0.55$) shows that the rows have norm close to 1. And density at 11 is very low, making $1$ a good separation point for two groups of row norms. The 17 rows of panels from top to down correspond to 17 $\alpha$ from 0 to 2. And the 6 columns from left to right correspond to $m=10,15,25,39,63,100$.
Here, we provide the heat map of $\phi_{1/2}$ at different $m$ as a function of $\alpha$ and $\gamma$ (Figure 13). The pattern is robust, suggesting weight decay is a good tool to change the degree of superposition regardless of data properties and model sizes.
Figure 13: Fraction of represented features as a function of $\gamma$ (x-axis) and $\alpha$ (y-axis). The 6 columns from left to right correspond to $m=10,15,25,39,63,100$. The colorbar is $\phi_{1/2}$ where purple means 11 and white means 0.
D.4 Sparsity does not affect scaling behaviors in our tests
Figure 14: Activation density does not affect scaling exponents in our tests. (a) Loss is roughly proportional to activation density given the degree of superposition ($m=63$, $n=1000$). (b) So, $E$ will only affect the coefficient but not the exponent when considering the power law with model dimension. We plot the evidence $\alpha_{m}\approx 1$ at strong superposition.
We studied the effect of the number of expected activated features or activation density $E$, which was set to $1$. By fixing data exponent $\alpha=1$, which will be shown to be relevant to natural language, we can scan different superposition degrees and activation densities. Since $p_{i}\leq 1$ is required, which is equivalent to $p_{1}\leq 1$, we have $E\leq\sum_{i=1}^{n}1/i^{\alpha}$, setting the upper bound for our scanning. We found that loss is approximately proportional to activation density $E$ (Figure 14a). This fact suggests that the power law exponent should not change, which we confirmed (Figure 14b). Under a controlled superposition degree, activation density linearly increases loss and thus does not affect the scaling exponents in our experiments.
@@ D.5 Figure 4
In Figure 4a, we plot the raw losses from small toy model experiments (hyperparameters in Appendix B.2). The theoretical value of loss is approximated by an integral $\int_{n\phi_{1/2}}^{n}1/i^{\alpha}\mathrm{d}i$, which is
$L=\left\{\begin{aligned} &\frac{\phi_{1/2}^{1-\alpha}-1}{1-n^{1-\alpha}}n^{1-\alpha},~\alpha\neq 1,\\ &-\frac{\ln\phi_{1/2}}{\ln n},~\alpha=1.\end{aligned}\right$. (14)
@@ D.6 Figure 5
Figure 16: Fraction of strongly represented features ($\|W_{i}\|_{2}>1$) at strong superposition ($\gamma=-1$) is around $\min\{1,m^{2}/2n\}$ (fast increasing dashed line), which is much larger than $m$ (slowly increasing dashed line). Colorbar means $\alpha$ as the main text.
For convenience, based on the observation that the vector norms are bimodal around $1$ in the strong superposition regime, we define strongly represented features as those that have $\|W_{i}\|_{2}>1$, which are more frequent and ETF-like, and weakly represented ones for those with $\|W_{i}\|_{2}<1$. We can quantify the fraction of strongly represented as
$\phi_{1}=|\{i:\|W_{i}\|_{2}>1\}|/n,$ (15)
which significantly exceeds $m/n$ and is around the ETF upper bound $m^{2}/2n$ (Figure 16). The group of vectors $\|W_{i}\|_{2}$ with norm greater than 1 or the strongly represented features then roughly agree with ETF properties: small variance, $1/m$ mean squared overlaps, and a limited number of vectors.
@@
Figure 5 studies the results from small toy models (Appendix B.2) focusing on the strong superposition regime. For the strongly represented fraction, $\phi_{1}$, we can directly compute based on the definition and the obtained weight matrices. We showed a row norm distribution at $m=15$, $\alpha=1$, and $\gamma=-0.78$ in Figure 5 panels a and b. Here, we provide more data to show that 1 is a natural separation point in norm to determine which are strongly represented and which are weakly represented (Figure 12).
@@
Once select the rows with norm greater than 1, we can calculate their mean and variance of squared overlaps based on normalized rows $W_{i}/\|W_{i}\|_{2}$ (Figure 5, c and d). We argue that after training, the vectors will be more similar to ETFs than to random initialization. This is studied via the variance of overlaps. ETFs, in theory, have zero variance. We find that the majority of the overlap variances are much smaller than the random initialization, especially when features have similar frequencies, which agrees with the expectation. The cases where the actual variance is greater than that of the random vectors have large $\alpha$, roughly correspond to the cases where $\alpha_{m}$ deviates from $1$ — ETF-like configuration no longer dominates. This is intuitive that when $\alpha$ is too large, the heterogeneity of overlaps will become large — it is better to let more frequent features occupy larger angle space. We argue that the large variance at large $\alpha$ does not mean the configuration tends to be random, but tends to be something more closely related to the frequency distribution of the features.
@@
Our explanations based on the strongly and weakly represented features capture the basic trend that when $\alpha$ is getting large, the more important features will have larger angle space and the loss decay will be more related to the data exponent. However, this theory is oversimplified, where the strongly represented all have small overlaps and the weakly represented all have large overlaps. The real situation may be more like the angle occupied by one feature decreases continuously as the frequency decreases. As suggested by Figure 5c, overlap variance within the strongly represented is greater when $\alpha$ is larger. To be more precise about the overlap distribution as well as the exponent $\alpha_{m}$ when $\alpha$ is large, we cannot use simple theoretical expectations like ETFs but have to solve the toy model.
Figure 17: The mean squared overlap over all the $W_{i}$ vectors in the strong superposition regime ($\gamma=-1$). Although the value may be higher than $1/m$ (the dashed line), the scaling is robustly $1/m$. Colorbar means $\alpha$ as the main text.
We also provide evidence that overlaps of all the vectors (Figure 17). We see that some of the mean square overlaps are larger than $1/m$ instead of being on the line $1/m$. However, all the mean values follow $1/m$ scaling even for large $\alpha$ cases where the vectors are no longer isotropic. We emphasize that for even frequencies and isotropic vectors, since squared overlaps scale as $1/m$, the loss should scale as $1/m$.
@@
Figure 18: Small toy models with $m$ from $50$ to $150$ (such that $m^{2}/2>n$) in the strong superposition regime yield similar $\alpha_{m}$ around 1 when $\alpha$ is small and a slightly smaller $\alpha_{m}$ (smaller than $2(\alpha-1)$) when $\alpha$ is large. We copied Figure 5e and made the points transparent for comparison. The non-transparent points are from small toy models with $m$ from $50$ to $150$. $\alpha_{m}=1$ is the horizontal line, $\alpha_{m}=2(\alpha-1)$ is the fast increasing dashed line, and $\alpha_{m}=\alpha-1$ is the slowly increasing dashed line. Error bars are standard errors. The colorbar encodes weight decay as in the main text.
@@ D.7 Figure 6
Figure 20: LLMs are in strong superposition based on the non-zero norms of the representation vectors. (a) OPT models. (b) Qwen2.5 models. (c) GPT-2 models. (d) Pythia models. The dots with error bars are mean values, and the error bar is the standard deviation of norms. Red small points are the maximum values, and dark small points are the minimum values. The mean norm value as a characteristic row norm does not depend on model size much as expected (Appendix A.2).
@@
[…]
@@
We study the relationship between model dimension $m$ and model size $N$ (number of parameters). For the four open-sourced models we analyzed [39, 40, 41, 42], we can see that $N\sim m^{3}$, especially when $m$ is large. If we fit the NN-$m$ relation by a power law while assuming a universal exponent but different coefficients depending on the model class, we obtain an exponent of $2.51$ (Figure 23a). For the Chinchilla model reported in [3], we find $N$ is also close to a power law with the model dimension, and the fitted exponent is $2.52\pm 0.03$ (Figure 23b).